Find the general integral of
(x^3-3xy^2)p+(y^3+3x^2y)q=2(x^2+y^2)z
1
Expert's answer
2020-12-17T07:18:02-0500

"\frac{dx}{x^3-3xy^2}=\frac{dy}{y^3+3x^2y}=\frac{dz}{2z(x^2+y^2)}"

"\frac{dy}{dx}=\frac{y^3+3x^2y}{x^3-3xy^2}"

This is a homogeneous differential equation.

Let "y\/x=v, y=xv, dy\/dx=v+x(dv\/dx)"

Then:

"v+x\frac{dv}{dx}=\frac{(xv)^3+3x^3v}{x^3-3x^3v^2}=\frac{v^3+3v}{1-3v^2}"

"x\frac{dv}{dx}=\frac{v^3+3v}{1-3v^2}-v=\frac{4v^3+2v}{1-3v^2}"

"\frac{1}{2}\intop\frac{1-3v^2}{v(2v^2+1)}dv=\intop \frac{dx}{x}"

"\frac{1-3v^2}{v(2v^2+1)}=\frac{A}{v}+\frac{Bv+C}{2v^2+1}"

"A(2v^2+1)+v(Bv+C)=1-3v^2"

"2A+B=-3"

"A=1"

"C=0"

"B=-3-2A=-5"

"\frac{1-3v^2}{v(2v^2+1)}=\frac{1}{v}-\frac{5v}{2v^2+1}"

"\intop\frac{v}{2v^2+1}dv=\frac{1}{2}\intop\frac{d(v^2)}{2v^2+1}dv=\frac{ln(2v^2+1)}{4}"


"\frac{lnv}{2}-\frac{5ln(2v^2+1)}{8}=lnx+lnc_1"

"ln(\frac{v^{1/2}}{(2v^2+1)^{5/8}})=ln(c_1x)"

"\frac{(y/x)^{1/2}}{x(2(y/x)^2+1)^{5/8}}=c_1"



"\frac{dx/x+dy/y+dz/z}{6x^2}=-\frac{3(dx/x-dy/y+dz/z)}{6y^2}=\frac{dz}{2z(x^2+y^2)}"

"\frac{dx/x+dy/y+dz/z-3(dx/x-dy/y+dz/z)}{6x^2+6y^2}=\frac{3dz/z}{6(x^2+y^2)}"

"dx\/x+dy\/y+dz\/z-3(dx\/x-dy\/y+dz\/z)=3dz\/z"

"4dy\/y-2dx\/x=5dz\/z"

"4lny-2lnx=5lnz+lnc_2"

"ln(y^4\/x^2)=ln(c_2z^5)"

"\frac{y^4}{x^2z^5}=c_2"


The general integral:

"\phi(\frac{(y/x)^{1/2}}{x(2(y/x)^2+1)^{5/8}}, \frac{y^4}{x^2z^5})=0"